Mark is playing a trivia game. On his first turn, he lost 100 points. On his second turn, he lost 55 points. On his third turn, he lost 65 points. Enter a sum that models the change to Mark's score after his first three turns.
1
Expert's answer
2020-05-01T16:54:46-0400

Solution:

1.  Mark starts with 0 points. Since he lost 100 points on his first turn, we can subtract:

0 - 100 = -100

2. From there, he lost another 55 points, meaning we subtract that from his first-turn total: -100 - 55 = -155

3. On his third turn, he lost another 65 points, so we subtract from his second-turn total:

-155 - 65 = -220

Answer:

Mark’s total points at the end of three turns is -220